

********************************************************************
UARS WEEKLY STATUS REPORT - 31 JULY 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE recently ended daylight only operations. During this limited operation
we did not take any stellar calibration measurements but successfully gathered
a full solar data set each day. Normal operations resumed on Thursday (98211).
***********
SUSIM
***********
SUSIM was turned off for most of this past week. It was successfully
turned on early Thursday. Since then, SUSIM has successfully gathered
its normal solar, data without apparent problems.
***********
PEM
***********
The PEM Zenith boom was activated Thursday, on July 30. PEM spectrometers
and magnetometer produce atmospheric energy deposition profiles from in situ
electron and proton measurements.
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of the OBC plan and its operation test procedures.
***********
HALOE
***********
HALOE was powered off at the end of day July 29, 1998. All engineering
parameters were nominal and the science data products continue to be excellent.
***********
MLS
***********
MLS continues to operate with incremental changes to scan software in order to
maximize visibility into scan mechanism and optimize science observations.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII is currently off, awaiting activation on August 8. The following
WINDII paper has just appeared in print:
Tidal influence on midlatitude airglow: Comparison of satellite and
ground-based observations with TIME-GCM predictions, G.G. Shepherd,
R.G. Roble, S.-P. Zhang, C. McLandress and R.H. Wiens, J. Geophys. Res.
103, 14741-14751, 1998.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 402 CD-ROM requests as of 31-July-1998.
One order was received and shipped.
The CDHF loaded May 1998 DAAC data into the DAAC Transfer Tables for
HALOE, HRDI, NMC, PEM, and UKMO.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
116 days HAL (4.80) PROD (L1) (1873R-815C-15F=1043L)
1 day HAL (4.80) PROD (L3) (1873R-559C=1314L)
138 days SOL (4.62) PROD (L2-L3) (2349R-835C=1514L)
1 day PEM (3.32) PROD (1R-1C=0L)
3 days SEN (5.0) PROD (4R-3C=1L)
0 days WIN (5.10) ENG. (1R-1F=1L)
The Level 3 Catalog Save containing UARS level 3A, 3L and 3T will cease on
08/01/98. If re-creation of the original file is needed in the future, the
algorithm will be reprocessed. CDHF users have been informed of this new policy.
CDHF pending software upgrades: IDL 5.1, OpenIngres, Fortran V6.
***********
FOT
***********
REPORTING PERIOD
This report covers 18 July 1998 (Orbit 37427, GMT Day 199) through 24 July
1998 (Orbit 37531, GMT Day 205).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 199/1011 37433 Scan Test
HALOE 200/0054 37442 On
HALOE 200/0235 37443 Low Power
HALOE 200/1026 37448 On
HRDI 202/0101 37472 On
HRDI 202/0252 37473 Low Power
HRDI 202/1035 37478 On
HRDI 204/2233 37515 Off
ACRIM 204/2233 37515 Off
SUSIM 204/2233 37515 Off
MLS 205/0113 37517 On
The beta angle ranged between 10.5 degrees and 34.6 degrees during this
report period. The beta angle is now increasing toward a maximum angle of
73.9 degrees on 05 August 1998 (DOY 217).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.8 to 1.2 degrees C. Additional battery performance data
for each battery on 18 July 1998 and 24 July 1998 are provided in Appendix
D.
The clock error ranged from 10.1 to -9.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 11
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 02
% ACRIM 01
% HALOE 03
% HRDI 05
% MLS 03
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 00
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/146 - 2287.496748
Transponder B: 98/147 - 2287.499000
GROUND SYSTEM OPERATION
No (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/204 show the observatory data loss to be
20 hours, 45 minutes, 33.748 seconds (no change since last report period).
This is a 0.0346 percent data loss which equals a 99.9654 percent data
capture for the mission.
Applications Processor (AP) ground system software testing has been
suspended waiting release of an operational pack containing version 10.11.
Received responses from two instrument teams - SOLSTICE and HALOE -
regarding inventory of the instrument microprocessor loads resident on the
Applications Processor (AP).
OTHER SIGNIFICANT EVENTS
The B/D thruster refill effort continues. Planning continues and the
engineering thruster tests goal is still 28 July 1998.
The magnetic field model has been calibrated and aligned using the UARS
Test and Training Simulator (UTTS). The TAM-controlled attitude
acquisition mode of the OBC ACS software was successfully tested. This
mode was added to the ACS software as part of last summerUs OBC software
V62 uplink (in conjunction with the epoch reset). The TAM-controlled
acquisition mode is intended to be used in place of the nominal ESAM-
controlled acquisition mode, in the event the update filter (UFLTR)-
controlled precision mode needs to be exited. This could happen in
response to an ACS anomaly (e.g., a gyro Rbit toggleS), or due a planned
circumstance (e.g., reconverging UFLTR with high-rate gyros prior to an
orbit adjust). Without the TAM-controlled acquisition mode, UFLTR
reconvergence and transition back to precision-pointing mode would have to
be performed
under ESA control. Due both to the unreliability of this mode and the
desire to preserve ESA life for safehold recovery procedures (for which
there is no alternative -- TAM control is not an option for safehold
recovery), it is highly preferable to use the TAM-controlled mode in place
of the ESA-controlled mode, if possible.
The testing was a complete success, with reported acquisition attitude
errors never exceeding 1.3 degrees, and within 1.0 degrees for nearly the
entire test period. Simulated ESA operation during the convergence showed
errors within 1.0 degrees during the entire test. UFLTR convergence ran to
completion in nearly the same amount of time it would have taken under ESA
control, indicating that the TAM-controlled mode is an operationally
acceptable mode to be used in place of ESA control.
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the UARS ground system phase-1 re-host effort to
the Lockheed Martin SCS-21 system. Work continued on the test script. The
Command Test (CMDTEST) procedure was successfully executed on the UARS
Applications Processor/UTTS on 23 July 1998. TDRSS input information for
the simulation was generated and provided to MSOCC System Engineer.
Several meetings/telecons with the SCS-21 personnel were supported to
clarify UARS requirements for UARS re-host.
MEETINGS
FOT met with Sidney Kirschner (CSC) of the CMS Year 2K test team on 23 July
1998 to discuss test procedures for the upcoming Year 2K CMS testing. In
addition, the FOT arranged for Sidney to receive some hands-on CMS training
over the next few weeks.
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 25 August
1998.
The UARS orbit adjust maneuver scheduled for 12 August 1998 has been
canceled and will be rescheduled at a later date.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 24 July 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
37431 199/0649 98-246 MLS NOOPS - prep for Scan Test
37433 199/1011 98-247 MLS Scan Test
37442 200/0054 98-248 HALOE Power ON
37443 200/0235 98-248 HALOE Low Power Contingency
37448 200/1026 98-249 HALOE Power ON
37472 202/0101 98-250 HRDI Power ON
37473 202/0252 98-251 HRDI Low Power Mode
37478 202/1035 98-252 HRDI Power ON
37478 202/1035 98-253 HRDI Htr Delay to 125 amp/min
37505 204/0627 98-254 OBC Clk Adj fr 0.00 to 20.6 msec/day
37515 204/2233 98-255 HRDI Power OFF
37515 204/2233 98-256 ACRIM Power OFF
37515 204/2233 98-257 SUSIM Power OFF
37517 205/0113 98-258 MLS Power ON
37530 205/2206 98-259 OBC Clk Adj fr 20.6 to 0.00msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 24 July 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIEUTEMP Y-LO WINDII EU Temperature 37427-37531
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 37430-32,62,76-82,
92-97,508-12, 522-31
SOSTMPRNTRN Y-LO Temperature Entrance 37431
HAH28VA R-LO Heater Motor +28v 37443-8
HASMIRRT Y-HI Secondary Mirror Temp 37444
HRPTSPBIMON Y-HI PTS Pulse Bus Current Monitor 37473,79,80,503
NBRECBI Y-HI NBTR B Current 37476,82
HRPSEQBIMON Y-HI PTS Quiet Bus Current Monitor 37482
NBBTAPEPOS R-LO NBTR B TAPE POSITION 37511
HRIFESURVTMP Y-HI HRDI IFE Survival Heater Temp. 37517-8
HRRADTEMP Y-HI HRDI IFE Radiator Temperature 37518
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 24 July 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
37505 204/06:32 (23 Jul 98) from 00.0 to 20.6 msec/day
37530 205/22:12 (24 Jul 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 24 July 1998
18 Jul 1998 - GMT Day 199 Beta = 10.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/ +0.0 18.2 0.000 0.0
2 +39.2/-22.4 +5.47/+4.51 +30.4/-13.6 25.1 1.023 22.0
3 +56.0/-11.2 +4.51/+2.92 +29.6/-14.4 25.1 1.020 21.5
24 Jul 1998 - GMT Day 205 Beta = 34.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.1 0.000 0.0
2 +16.8/ -33.6 +6.76/+4.51 +29.6/-16.0 24.3 1.018 23.0
3 +33.6/ -33.6 +5.47/+3.55 +28.8/-16.0 24.2 1.021 22.4
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 24 July 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-044 FOT See list July 1998 14170 Generic Late Acquisition